You just took out a $350,000, 25-year mortgage at a 5-year rate of 5% from Royal Bank. You will make monthly payments.

Assume that the mortgage rate remains at 5% for the remaining time of the mortgage. What is your monthly mortgage payment? What are the principal and interest payments for the 200th monthly payment?
Answer #1

1.
Monthly payment=PMT(5%/12,12*25,-350000)=$2,046.07

2.
Interest payment=FV(5%/12,199,2046.07,-350000)*5%/12=$701.64
Principal payment=2046.07-701.64=$1,344.43